OpenStudy (anonymous):

simplify x^2-9 over x^2-6x+9

OpenStudy (amistre64):

notice that you have a difference of squares over a perfect square

OpenStudy (amistre64):

factoring or long division might work

OpenStudy (amistre64):

1 + (6x-18)/(x^2-6x+9) ------------------ x^2-6x+9 | x^2 -9 -(x^2-6x+9) ------------- 6x-18 1 + 6(x-3)/(x^2-6x+9) gotta factor either way .... so id stick to factoring

OpenStudy (amistre64):

x^2-9 = (x+3)(x-3) --------- x^2-6x+9 = (x-3)(x-3) cancel out like terms

OpenStudy (anonymous):

thank you.